SITE 1072
(Proposed Site MAT-9B-1)

Hole 1072A

The ship was moved in DP mode 1.9 nmi along the trend of site-specific seismic lines bearing ~125° from Site 1071 to the dGPS coordinates for proposed Site MAT-9B-1. Observations of the tethered beacon vs. the constrained beacon at Site 1071 suggested that the beacon signal problem was related to beacon offset distance and not current disturbance of beacon attitude. Two-meter tethers were used again because they were successful at Site 1071; historically, shorter tethers have caused the beacons to collide with the weights on impact. The ship was positioned with the moonpool over Hole 1072A for both beacon drops. The first beacon (15.0 kHz) was dropped at 0723 hr on 29 June and moved 10 m at 161° before settling. A second beacon (17.0 kHz) was dropped at 0750 hr on 29 June and moved 23 m at 187°. A TV and sonar survey was conducted on a 6-m square pattern centered on the site coordinates, with no bottom obstructions observed. One of the beacons experienced a signal loss when the hole was spudded (about 3 m away), and periodically thereafter during wireline operations, which suggested that, ideally, beacons should be offset about 7 m from the hole to avoid acoustic interference. The one working dGPS unit was more stable than regular GPS; nevertheless, it experienced periods when it drifted despite efforts to stabilize it. Shore-based assistance was sought to get the other dGPS unit to work. The occasional drift in dGPS (±4 m ) may have exacerbated apparent beacon drift.

The RCB coring system was selected over APC/XCB to determine if the anticipated increase in ROP and smaller bit diameter might produce a more stable hole and achieve better recovery than was achieved at Site 1071. The seafloor was tagged at 98.2 mbsl (109.5 mbrf), and Hole 1072A was spudded at 1400 hr on 29 June. RCB Cores 1R through 51R were recovered from 0 to 282.8 mbsf. Half-cores (5-m advances) were taken to increase recovery. Mud sweeps were required to keep the hole clean after nearly every core (and sometimes every half-core) over the entire interval, using 920 bbl of sepiolite mud in Hole 1072A (four times the normal usage). After Core 51R, increasing pump pressures and torque indicated possible hole sloughing, but mud sweeps cleared the problem. Coring continued with RCB Cores 52R through 55R from 282.8 to 301.8 mbsf. Again, torque and pressure indicated continuing hole problems; therefore, a wiper trip was made to 229.6 mbsf. Core 56R was recovered from 301.8 to 306.8 mbsf, but hole problems persisted despite additional mud sweeps.

Another wiper trip was made to 214 mbsf, and the hole was displaced with sepiolite mud in an effort to stabilize the sands for logging. The pipe was pulled to the ship with heavy backflow to 91 mbsf. The seafloor was then surveyed with the VIT-TV and sonar; a 4-m-diameter slightly irregular crater at the seafloor was observed. The closer beacon was noted to be on the edge of this crater. The hole appeared to be back-flowing water and sand, possibly from being overpressured during drilling. A wiper trip was made to 296 mbsf, and 20 m of fill was tagged. The hole was displaced with sepiolite mud again, and the RCB bit was pulled to the ship to prepare for logging.

The open hole was reentered with sonar because the crater was obscured by what appeared to be flowing sand and water. The bit was stopped at 63 mbsf for logging. A triple combo (DITE/HLDT/APS [slim-hole lithodensity tool/accelerator porosity sonde]) log was run to 300 mbsf (9 m off bottom) in 2.75 hr. While preparing for the next logging run, the drill pipe began to stand up out of the elevators, indicating that the pipe was stuck. An unsuccessful effort was made to free the pipe by working it in gradually increasing increments from 20,000 to 200,000 lb overpull and by circulating through the hose at pump pressures from 500 to 2500 psi at 300 1000 gpm. It was not possible to pick up the top drive or break the connection. A string shot was fired below the bit in hopes that the sand bridge could be knocked loose; however, the pipe remained stuck. The pipe was severed with an explosive charge at the top of the tapered drill collar at 23.5 mbsf and was pulled free with 100,000 lb overpull. The pipe cleared the rotary table at 1500 hr on 2 July, ending Hole 1072A.

Hole 1072B
Subsequent discussions concluded that coring delays (from wireline operations) and hole disturbance from the additional circulation required to clean Hole 1072A (so circulation could be stopped to retrieve core barrels) could have aggravated hole stability in sandy intervals. We concluded that a smaller dedicated logging hole could be drilled and logged quickly, with better hole conditions. Therefore, the ship was moved in DP mode 20 m southeast along a seismic line bearing 125° from Hole 1072A. Observations of the beacon farther from Hole 1072A proved that the signal was not adequate for positioning at the new hole; therefore, the beacon was released. The ship was positioned with the moonpool over the intended spud point, and a new beacon (15.0 kHz, 193 dB) was dropped at 1530 hr on 2 July; however, the beacon moved 10 m at 304°, apparently due to current. The initial operational plan was to drill a LWD hole to 300 mbsf, and possibly also run wireline logs; however, a decision was made to drill the hole for wireline logs first, then further evaluate formation stability before attempting a LWD run. Hole 1072B was spudded at 2340 hr on 2 July and drilled to 306.8 mbsf at 20.1 m/hr with sepiolite mud sweeps every other connection (19 m). A wiper trip was made to 71 mbsf, with 10,000 lb maximum overpull and drag and flow-back pressure to 90 mbsf and 10 m fill. A second wiper trip was made to 250 mbsf, and the hole was filled with sepiolite again.

The persistent hole problems may have been due in part to weak and unstable silty sands at 30-40 mbsf, which sloughed and enlarged the hole diameter and/or fractures and resulted in lost circulation. Both effects eventually resulted in reduced annular fluid velocity, which caused cuttings to collect in the upper hole rather than circulate out to the seafloor. The cuttings eventually formed a bridge, impeded vertical flow, increased annular pressure (and lost circulation), and packed-off the hole, causing the pipe to stick.

The seafloor was surveyed with the VIT-TV and sonar on the trip for the logging bit. The open hole was obscured by a boiling cloud (suggesting possible flow-back from drilling operations), but it was reentered in 1 hr using sonar. The bit was positioned at 45 mbsf, because lateral motion there would be confined by unusually stiff clays near the seafloor. The DITE/Sonic/NGT tool was run, but it would not pass 90 mbsf. A wiper trip was made from 30 to 102 mbsf with circulation and light reaming. The hole was displaced with sepiolite mud, and the pipe was positioned again at 45 mbsf. The DITE/Sonic/NGT and FMS/NGT tools were run successfully to 295.5 mbsf (11.3 m off bottom) in 4 hr. A vertical seismic profile (VSP) log was run to the same depth with Schlumberger's well seismic tool (WST) tool in 3 hr. The reentry/logging bit cleared the rotary at 2100 hr on 4 July.

The RCB BHA was rerun, and Hole 1072B was reentered. The bit took weight at 17 mbsf, indicating persistent closure of the upper hole from 20 to 60 mbsf. The hole was reamed through five tight spots and 8.5 m fill to 306.8 mbsf TD. RCB Cores 1072B-1R through 6R were cut from 306.8 to 358.6 mbsf; however, coring was terminated after penetrating nearly 100 m of presumed unconsolidated sand with negligible recovery. Continued penetration of the sands may have endangered another BHA without any significant scientific return, so it was judged prudent by all parties to cease operations at Hole 1072B. The pipe was pulled at 1610 hr on 5 July, ending Hole 1072B.

Hole 1072C
The successful drilling and logging operations conducted in Hole 1072B indicated that a nearly identical hole could safely be drilled using Schlumberger-Anadrill LWD tools. The ship was moved 30 m at a 125° heading from Hole 1072B to assure the best possible hole conditions and avoid any disturbance from lost circulation at the previously drilled holes. A beacon (15.0 kHz, 193 dB) was deployed through the moonpool at 1650 hr on 5 July, which moved 4.3 m at 310°. Another beacon (15.0 kHz, 193 dB) was deployed through the moonpool at 1700 hr and moved 21 m at 265°. The ship was moved another 5 m at a 125° heading to maintain offset distance from the first beacon.

The initial LWD plan called for drilling to 125 mbsf without the jars (to avoid flexing the jars above the seafloor), pulling out of the hole and picking up the jars, then reentering and wiping through the upper hole before continuing. The bit tagged bottom at 99.5 mbsl (111.0 mbrf), and Hole 1072C was spudded at 2030 hr on 5 July and drilled to 106.9 mbsf in 6 hr. The ROP was controlled to 25 m/hr. The bit was pulled above the seafloor to add the jars; an overpull of 30,000 lb up to 91 mbsf and 10K lb up to 29 mbsf was noted, indicating that the upper hole was closing in or packing off. The jars were added to the BHA, and the VIT was deployed for reentry; however, an electrical short was noted in the TV cable. The VIT repair was estimated to take 18 hr; therefore, a decision was made to continue LWD logging, either with a blind reentry into Hole 1072C or by spudding a new hole. The seafloor crater for Hole 1072C was located by offsetting the ship a few feet and feeling for bottom with the bit, but efforts to make a blind reentry were not successful. Hole 1072C officially ended at 0335 hr on 6 July when the bit cleared the seafloor.

Hole 1072D
The ship was moved 14.6 m at a heading of 305° from Hole 1072C, back toward Hole 1072B. The LWD BHA was reconfigured by taking out two drill collars, so the jars could be run at spud-in and no reentry was required. The seafloor was tagged at 99.5 mbsl (111.0 mbrf), and Hole 1072D was spudded at 1030 hr on 6 July. The LWD hole was drilled from 0 to 110.4 m. Previous hole problems dictated that extra precautions be taken with the LWD tools, jars, and BHA. Soft clays from 0 to 30 mbsf had been squeezing into the annulus and packing-off previous holes; therefore, a precautionary wiper trip was made with the top drive to 32.8 mbsf, and the BHA was run back to bottom with 10,000 lb maximum drag and 4 m of fill. Drilling continued in the LWD hole from 110.4 to 233.4 mbsf. A second precautionary wiper trip was made with the top drive to 32.8 mbsf with 10,000 lb maximum overpull, and the BHA was run back to bottom with 5,000 lb maximum drag and 18 m fill. Drilling continued in the LWD hole from 233.4 to 356.0 mbsf. The LWD BHA was pulled and successfully cleared the rotary at 1825 hr on 7 July, officially ending Hole 1072D.

To 174A Operations - Site 1073

To 174A Table of Contents

ODP Publications

ODP Homepage